OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = -7 x - 3y + 4z = -14 x - 2y - 3z = -11

OpenStudy (anonymous):

I dont know how to go about this one

OpenStudy (shamil98):

Multiply the bottom two equations by -1 and then add the equations (this will get rid of the x and z variables) to solve for y.

OpenStudy (shamil98):

Got it from there?

OpenStudy (anonymous):

Is that all i have to do?

OpenStudy (shamil98):

pretty much but put it in your own words, as it says "include the solution to one of the variables"

OpenStudy (shamil98):

that would be the process of elimination i think, i forget the terminology used.

OpenStudy (anonymous):

-x + 3y - 4z = +14 -x + 2y + 3z = +11 y-z=3?

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

I shill have two variables

OpenStudy (shamil98):

no, the question says to find a solution to ONE of the variables.

OpenStudy (anonymous):

Oh so y-z=3 is my answer

OpenStudy (anonymous):

wait 14+11 is 25 im so dumb

OpenStudy (anonymous):

i was seeing 4-1 for some reason i have no idea why

OpenStudy (shamil98):

no.. 2x + y + z = -7 x - 3y + 4z = -14 x - 2y - 3z = -11 multiply the bottom two by -1 you get 2x + y + z = -7 -x +3y - 4z = 14 -x +2y +3z = 11 add them all x and z cancel out so you're left with 6y = 18 y = 3

OpenStudy (anonymous):

i was also seeing 3-2 lolol i need to take a break ive been doing this for too long lol

OpenStudy (anonymous):

thank you!

OpenStudy (shamil98):

yeah np